# SpikingBrain: Spiking Brain-inspired Large Models

### v2

@@ 5.1 Downstream Performance
Table 3: Performance Evaluation of SpikingBrain Chat Models. All models are tested with the vLLM framework and evaluated using a generation-based method. Except for Qwen2.5, the other baselines are trained on limited Chinese data, resulting in clear disadvantages on CMMLU and C-Eval. For QuALITY and IFEval, we report results from the non-CoT model (after SFT stage 2) to avoid chain-of-thought interference.
[…]
SpikingBrain-7B
[…]
SpikingBrain-76B
@@
MMLU (hendrycks2020mmlu, ) 65.57 73.71 68.69 75.17 71.03
CMMLU (li2023cmmlu, ) 68.76 77.41 55.17 79.14 51.03
HS (zellers2019hellaswag, ) 68.95 86.63 76.80 85.39 75.63
Ceval (huang2023ceval, ) 69.07 76.32 55.01 77.93 50.88
NQ (kwiatkowski2019nq, ) 21.47 21.55 30.97 17.67 28.48
TrQ (joshi2017triviaqa, ) 57.03 55.13 65.78 55.72 71.00
QuALITY (pang-etal-2022-quality, ) 60.12 69.56 66.25 73.63 51.34
IFEval (zhou2023instruction, ) 42.70 49.72 73.01 73.20 48.06
5.2 Long-context Efficiency
Beyond task performance, we also place particular emphasis on the efficiency of our models in long-sequence inference. Benefiting from the integration of (hybrid) linear attention with the MoE mechanism, both SpikingBrain models exhibit substantial inference speedups. Since the 7B model is more compact and shows stronger efficiency gains, we adapt it to the HuggingFace framework with multi-device sequence-parallel inference (ZeCO + p2p), successfully supporting efficient prefill up to 4M tokens. As shown in Table 4 and Figure 6, at an input length of 1M tokens, the SpikingBrain-7B model achieves a 26.5× speedup in TTFT (Time to First Token) compared with the Qwen2.5 baseline using full attention and all-to-all communication. Due to resource limitations and the number of attention heads, Qwen2.5 is difficult to evaluate at 4M tokens. Based on the fitted scaling curve, we conservatively estimate a speedup of over 100×. Moreover, as sequence length and GPU count increase proportionally, the time overhead of our model remains nearly constant. These results clearly demonstrate the scalability and superiority of our approach in ultra-long-sequence scenarios, providing strong support for the deployment of more efficient large language models in practical applications.
@@ 5.2 Long-context Efficiency
Sequence length GPU count SpikingBrain-7B Qwen2.5-7B
256k 8 1015 7419
512k 16 1037 14398
1M 32 1054 27929
2M 64 1070
4M 128 1073
For long-sequence training, we further measure Throughput per GPU Second (TGS) on SpikingBrain-7B and the Qwen2.5 baseline under the sequence parallelism framework. The experiments show a 5.36× throughput advantage for our 7B model at a sequence length of 128k, consistent with the TTFT improvements observed during inference.
@@
5.3 CPU-side Inference
In this section, we demonstrate the deployment of compressed 1B-scale models, including the SpikingBrain linear model with SWA and GLA, as well as Llama3.2 (dubey2024llama, ), on a CPU-based mobile inference framework. The deployment leverages llama.cpp as the inference backend. As illustrated in Figure 7, the workflow consists of four main steps:
@@ 5.3 CPU-side Inference
  • Graph and operator optimization. Following the forward propagation flow, we implement key operators including attention, rotary position embedding (RoPE (Su2021RoFormerET, )), and residual connections. Two major optimizations are applied: i) Operator fusion: combining matrix multiplication, bias addition, and nonlinear activation into a single kernel to reduce intermediate storage and memory access. ii) Graph-level optimization: decomposing and reorganizing operations such as RMSNorm, Softmax, and RoPE to enable fusion with matrix multiplication or low-level kernels, minimizing redundant computation.
@@
5.4 Performance on the MetaX Cluster
The performance of the MetaX C550 GPU cluster is evaluated from two perspectives: training efficiency and system stability. For efficiency, we report model FLOPs utilization (MFU) and tokens per second per GPU (TGS). During training, the cluster demonstrates consistently competitive performance. For the SpikingBrain-7B model, we achieve a TGS of 1558 and an MFU of 23.4% (8-way DP, 4-way PP, PP micro-batch size 2, global batch size 512), reflecting high computational efficiency and effective resource utilization.
Overall, these results highlight the capability of the MetaX GPU cluster to efficiently and reliably support large-scale long-duration training tasks, including the stable training of models with tens of billions of parameters.
5.5 Analysis of Spiking Scheme
Scheme deployment and accuracy maintenance
Winogrande Arc-e
Arc-c
HellaSwag
PIQA MMLU CMMLU Avg.
SpikingBrain-7B -
W8ASpike
0.6895 0.7861 0.5410 0.6758 0.7979 0.6546 0.6677 0.6875
SpikingBrain -76B-
W8ASpike
0.7148 0.7949 0.5371 0.6863 0.8004 0.7081 0.7512 0.7133
Spike distribution characteristics
To evaluate the sparsity and efficiency of the proposed spiking scheme, we analyze the statistical firing patterns under the bitwise–ternary coding method. We collect the absolute values of input spike integers across all linear projection layers. The detailed distribution of spike counts is illustrated in Figure 9. Take the SpikingBrain-7B model for example, results show that 94.1% of values fall within [0,7], and only about 1% exceed 16. After expansion into bitwise–ternary spikes, the average number of spikes fired per channel is only 1.13, while 18.4% of channels do not fire any spikes at all. Since 94.1% of values are within 7 (i.e., within the INT3 range), we further analyze the firing rate under a step size of 3. By filling inactive bits within the 3-step window with 0, the overall sparsity reaches approximately 69.15%.
Spike visualization interface
Energy Efficiency with Asynchronous Hardware
For computation, we estimate and compare the energy consumption of MAC (multiply–accumulate) operations under different paradigms, based on published hardware energy consumption data at 45nm technology (Speck, ):
    $E=\text{Average Spikes}\times E_{\text{INT8 Add}}.$ (29)
6 Conclusion
This work provides a comprehensive demonstration of efficient brain-inspired large model training on the MetaX GPU cluster. By integrating several key techniques—including novel (hybrid) linear architectures beyond Transformers, sparse MoE design, lightweight conversion training, and adaptive-threshold sparse spiking activation—we validate and implement a practical development pipeline for spiking-based large models on a non-NVIDIA cluster with hundreds of GPUs.
Acknowledgements
Author contributions statement
Yuqi Pan and Yuhong Chou designed the model architectures. Siyu Ding proposed the adaptive threshold spiking neuron and conducted spiking-based training. Han Xu and Xuerui Qiu developed three spike coding techniques. Yuqi Pan, Yupeng Feng, and Zehao Liu carried out the framework adaptation on MetaX and model training. Jinghao Zhuang, Yupeng Feng, Zehao Liu, Yuqi Pan, Bohan Sun, and Yuhong Chou performed model evaluation. Anjie Hu and Anlin Deng contributed to code adaptation and refinement. Peng Zhou performed CPU adaptation and model testing. Yuqi Pan, Guoqi Li, Man Yao, Jibin Wu, Jian Yang, Guoliang Sun, and Bo Xu contributed to the discussion and analysis of experimental results. The project was supervised by Bo Xu and Guoqi Li.

### v3

@@ A.3 Evaluation of Instruct Models
Table 4: Performance Evaluation of SpikingBrain Instruct Models. All models are tested with the vLLM framework and evaluated using a generation-based method. For QuALITY and IFEval, we report results from the non-CoT model (after SFT stage 2) to avoid chain-of-thought interference.
@@
MMLU 65.57 73.71 68.69 75.17 71.03
CMMLU 68.76 77.41 55.17 79.14 51.03
HS 68.95 86.63 76.80 85.39 75.63
Ceval 69.07 76.32 55.01 77.93 50.88
NQ 21.47 21.55 30.97 17.67 28.48
TrQ 57.03 55.13 65.78 55.72 71.00
QuALITY 60.12 69.56 66.25 73.63 51.34
IFEval 42.70 49.72 73.01 73.20 48.06
A.4 Evaluation on LongBench
To assess real-world long-context understanding, we evaluate SpikingBrain-76B (after stage‑2 SFT) on six LongBench tasks (Bai et al., 2024), including HotpotQA, NarrativeQA, TriviaQA, TREC, MultiNews and Qasper. Given the baseline’s maximum positional embeddings, we truncate prompts to 32k. As shown in Table 5, our model is competitive with other hybrid and Transformer baselines while using only four global softmax attention layers.
Table 5: Performance Evaluation of SpikingBrain-76B on six LongBench tasks.
[…]
Params 12B/76B 12B/52B 7B 13B/47B
Complexity Type Hybrid Hybrid Quadratic Quadratic
Benchmarks
HotpotQA 50.00 11.38 57.62 36.42
NarrativeQA 23.92 3.73 27.96 7.44
TriviaQA 90.84 92.63 85.92 85.97
TREC 67.00 77.00 68.50 19.25
MultiNews 25.89 8.44 23.79 26.52
Qasper 39.59 10.67 45.13 28.37
A.5 Long-context Efficiency
@@
Sequence Length 256k 512k 1M 2M 4M
GPU Count 8 16 32 64 128
[…]
SpikingBrain-7B
1015 1037 1054 1070 1073
[…]
@@
A.6 CPU-side Inference
We demonstrate the deployment of compressed 1B-scale models, including the SpikingBrain linear model with SWA and GLA, as well as Llama3.2 (Dubey et al., 2024), on a CPU-based mobile inference framework. The deployment leverages llama.cpp as the inference backend. As illustrated in Figure 8, the workflow consists of four main steps:
@@ A.6 CPU-side Inference
  • Graph and operator optimization. Following the forward propagation flow, we implement key operators including attention, rotary position embedding (RoPE (Su et al., 2021)), and residual connections. Two major optimizations are applied: i) Operator fusion: combining matrix multiplication, bias addition, and nonlinear activation into a single kernel to reduce intermediate storage and memory access. ii) Graph-level optimization: decomposing and reorganizing operations such as RMSNorm, Softmax, and RoPE to enable fusion with matrix multiplication or low-level kernels, minimizing redundant computation.
@@
A.7 Spike Visualization Interface
We provide a unified visualization interface to examine the spiking activity of neurons under different coding strategies. As shown in Figure 10, we compare Binary, Ternary, and Bitwise Spike Coding using two-dimensional time–neuron firing maps. This tool provides an intuitive means for analyzing spiking behaviors across different spike coding strategies.